The coefficient of x in the Maclaurin's expansion of sinx sink is


1
Expert's answer
2021-05-10T17:10:15-0400

Given, "f(x)=sinx sink"


so

"f'(x)=cosx sink\Rightarrow f'(0)=cos0 sink=sink
\\
f''(x)=-sin xsink\Rightarrow f''(0)=0
\\
f'''(x)=-cosxsink\Rightarrow f'''(0)=-sink"


Maclaurins expansion is given by-


"f(x)=f(0)+xf'(0)+\dfrac{x^2}{2!}f''(0)+\dfrac{x^3}{3!} f'''(0)...

\\[9pt]

f(x)=0+x(sink)+\dfrac{x^2}{2!}\times 0+\dfrac{x^3}{3!} (-sink)"


The coefficient of x in the expansion is sink.